OpenStudy (anonymous):

an airline has 82% of its flights depart on.schedule. it has 65% of its flights depart and arrive on schedule. find the probability that a flight departs on schedule also arrives on schedule. round the answer go two decimal places. (1 point)

OpenStudy (kropot72):

Let D be the event 'departs on schedule' and let A be the event 'arrives on schedule'. \[\large P(D)=0.82\] \[\large P(D \cap A)=0.65\] \[\large P(A)=\frac{P(D \cap A)}{P(D)}=\frac{0.65}{0.82}\] \[\large P(D \cup A)=P(D)+P(A)-P(D \cap A)\] The probability that a flight that departs on schedule also arrives on schedule is therefore: \[\large P(D \cup A)=0.82+\frac{0.65}{0.82}-0.65=you\ can\ calculate\]